OpenStudy (anonymous):

The length of a rectangle is 2 m less than twice the width. If the perimeter is 11 m, find the length and width

OpenStudy (anonymous):

fine I give up on this question

OpenStudy (anonymous):

damed

OpenStudy (wolf1728):

L + W = 11 L-2 = 2W

OpenStudy (anonymous):

ok ?

OpenStudy (wolf1728):

No - not done yet

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what variable do u want the width to be?

OpenStudy (anonymous):

don't know

OpenStudy (anonymous):

name any alphabet i will see that u understand how to tackle this pro

OpenStudy (anonymous):

i mean problem

OpenStudy (anonymous):

x

OpenStudy (anonymous):

fine  got it

OpenStudy (anonymous):

May I join friends :)

OpenStudy (anonymous):

YES SURE !

OpenStudy (anonymous):

yes by all means

OpenStudy (anonymous):

k

OpenStudy (anonymous):

don't get it

OpenStudy (anonymous):

as per the question the length is 2m less than twice the breadth twice the breadth gives = 2x 2m less gives 2x-2 got it?

OpenStudy (anonymous):

breadth is the same as width

OpenStudy (anonymous):

so what is the equation

OpenStudy (wolf1728):

I think my second equation is wrong. Working out the answer by trial and error Think it is Length = 6.66666 Width = 4.3333333

OpenStudy (anonymous):

@darkknight321: Can u write for me the formula to find the rigtangle around :) PLEASE :)

OpenStudy (anonymous):

 perimeter of any rectangle on this earth=2(length + breadth) length=2x-2 breadth=x perimeter=11 11=2(2x-2+x) 11/2=2x-2+x<--------------------taking the 2 on otherside of the equation 5.5=3x-2 solve for x(width) and then reply!